A taho vendor carries a 1.5 m long light plank over his shoulder. At the ends of the plank are two buckets weighing 40 N and 60 N respectively. (a) Find the value of force F exerted by his shoulder. Neglect the weight of the plank. ( b) Where should he support the plank for it to be balanced horizontally ( refer to FIG 1) 


1
Expert's answer
2022-02-01T09:57:05-0500

(a) Find the value of force F exerted by his shoulder:


"F=W_1+W_2=100\text{ N}."


(b) The point of equilibrium can be found from moments:


"W_1x=W_2(l-x),\\\space\\
x=\dfrac{lW_2}{W_2+W_1}=1.13\text{ m}"

from W1 (from the 40N-weight).